Find the angle between u (3.2, 7.2) and (0.3, 3.1) . Round to the nearest tenth. V
Answer #1

By using dot product of both vector

Go= u Criven ū = 3.20 +7.2} , 0 = 0.3ê +3.2 } ůů= lullulGO {o= angle between (32C+723) (0.32+117) lulle Je-272+1728]|0:38+131